lindsaylynn1120:

helpppppp matthhhhh agiannn An office manager orders one calculator or one calendar for each of the office's 80 employees. Each calculator costs $10, and each calendar costs $15. The entire order totaled $1,000. Part A: Write the system of equations that models this scenario. (5 points) Part B: Use substitution method or elimination method to determine the number of calculators and calendars ordered. Show all necessary steps. (5 points)

Marlin24k:

Did i get the last one correct? Just asking.

lindsaylynn1120:

yeah

Marlin24k:

Okay i'll help with the equation. 80x+10+15=1,000

lindsaylynn1120:

yeah ik that i nee d part b

Marlin24k:

Oh okay hold on.

lindsaylynn1120:

srry i should have been more supicific

Marlin24k:

It's fine i get it.

lindsaylynn1120:

alright

Marlin24k:

So do i just solve the equation?

Marlin24k:

Or what?

KyledaGreat:

Part A: Let c be the number of calculators ordered. Let d be the number of calendars ordered. The system of equations that models this scenario is: c + d = 80 (equation 1) 10c + 15d = 1000 (equation 2) Part B: Using substitution method: From equation 1, we have: c = 80 - d Substituting this into equation 2, we get: 10(80 - d) + 15d = 1000 Expanding and simplifying, we get: 800 - 10d + 15d = 1000 5d = 200 d = 40 Substituting d = 40 into equation 1, we get: c + 40 = 80 c = 40 Therefore, the office manager ordered 40 calculators and 40 calendars.

Marlin24k:

Only needed part "B" lmao.